OpenStudy (anonymous):

5g=3(h-1) what is h ?

OpenStudy (amistre64):

we would have to solve for h in terms of g

OpenStudy (amistre64):

divide out the 3 nd get rid of the 1

OpenStudy (anonymous):

THANKS , so what would the final answer be ?

OpenStudy (amistre64):

ill show you mine if you show me yours :) take a stab at it and lets compare ..

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

5g/3+1 ?

OpenStudy (amistre64):

5g=3(h-1) ; /3 5g/3 = h-1 ; +1 (5g/3) +1 = h .... thats what I get too, but the format might be off depending on whos doing the grading

OpenStudy (amistre64):

(5g+3)/3 = h is also acceptable

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

for every serving off potatoes Manny boils 1 1/2 potatoes . How many servings of potatoes can be made from 15 potatoes ?

OpenStudy (anonymous):

please help with this